OpenStudy (anonymous):

Holy crap, Someone explain this to me! Solve this system of equations. Please write answer as an ordered pair. -2x+6y=6 -7x+8y=-5

OpenStudy (jdoe0001):

can you solve the 1st for "y" ? what would you get?

OpenStudy (anonymous):

Problem is, I don't know where to start, everyone who did try to explain it.. well. i got more lost.

OpenStudy (jdoe0001):

well, just start by solving the 1st equation by "y" ? :)

OpenStudy (anonymous):

um.. I'll try..

OpenStudy (anonymous):

can someone please explain to me too???!!!

OpenStudy (jdoe0001):

heehe, sure =)

OpenStudy (anonymous):

I got 3..

OpenStudy (anonymous):

Thats not right..

OpenStudy (jdoe0001):

3? y = 3? whatever happen to the "x"? it went "missing in action" =)

OpenStudy (anonymous):

Hahaha, I don't know hahaha

OpenStudy (anonymous):

wolfram alpha! =]]

OpenStudy (anonymous):

Y= 3 x

OpenStudy (anonymous):

something, idk what I'm saying xD

OpenStudy (anonymous):

-2x+6y= 6 in slope intercept form is what?

OpenStudy (anonymous):

Hell, I dunno.. I'm so lost!!!!

OpenStudy (dbzfan836):

Ill explain it all

OpenStudy (anonymous):

y=1/3x+1 right? just add 2x to the other side and divide by 6. That's what slope intercept form is hence its name

OpenStudy (dbzfan836):

y=3 means x can equal anything, but y will ALWAYS equal 3

OpenStudy (anonymous):

now substitute y=1/3x+1 to the y in -7x+8y=-5

OpenStudy (anonymous):

1) Solve for y 2) plug your answer into one of the equations to get x 3)use your answer for and plug it one of the equations to find y

OpenStudy (dbzfan836):

-2x+6y=6 now, to do this, start off with graph paper, and make a graph

OpenStudy (texaschic101):

I would divide the first equation by a -2 and then solve for x. Then sub that into the other equation to find y. (-2x + 6y = 6) / -2 x - 3y = -3 x = 3y - 3 now sub that into the other equation and solve for y. Once you have y, sub that into either equation and find x.

OpenStudy (jdoe0001):

\(\bf 2x+6y=6\implies 6y=6-2x\implies y=\cfrac{6-2x}{6} \\ \quad \\ \implies y=\cancel{ \cfrac{6}{6} }-\cfrac{\cancel{ 2 }x}{\cancel{ 6 }}\implies ?\)

OpenStudy (dbzfan836):

Ok, listen up, @texaschic101 @jdoe0001 @lacrosseplayer22 @SaiVal

OpenStudy (dbzfan836):

I have an A+ in this thing. let me explain

OpenStudy (anonymous):

My god... This is another language...

OpenStudy (jdoe0001):

w0t? =)

OpenStudy (anonymous):

@Dbzfan836 Im done with Algebra in fact I was put into Adv. Algebra cuz I know it so well. no need to brag

OpenStudy (dbzfan836):

-2x+6y=6 say y is 0 here. what would x be to make the equation true? do the same to x -7x+8y=-5

OpenStudy (dbzfan836):

repeat, and record it

OpenStudy (jdoe0001):

hmmm there's a - in front of the 2... didn't notice it =)

OpenStudy (jdoe0001):

so @Keridavis can you solve for "y" on the 1st one? have you covered linear simplification already?

OpenStudy (dbzfan836):

-2x+6y=6 y=0 -2x+6*0=6 -2x+6y=6 x=0 -2*0+6y=6

OpenStudy (dbzfan836):

-2x+6y=6 y=0 -2x+6*0=6 [-3,0] -2x+6y=6 x=0 -2*0+6y=6 [0,1]

OpenStudy (dbzfan836):

make a line using those dots

OpenStudy (anonymous):

ugh... forget it.. I'm gonna go guys.. I'm sorry.

OpenStudy (dbzfan836):

@keridavis dont leave

OpenStudy (anonymous):

Why can't I leave?

OpenStudy (dbzfan836):

-7x+8y=-5 y=0 -7x+0=-5 -7x+8y=-5 x=0 0+8y=-5

OpenStudy (dbzfan836):

because were almost done, @Keridavis

OpenStudy (anonymous):

I don't understand anything you guys are saying.. Honestly, your wasting your time..

OpenStudy (dbzfan836):

ill just give u the answer

OpenStudy (dbzfan836):

one sec

OpenStudy (texaschic101):

reduce the first one...divide everything by -2, you will get x = 3y - 3. now sub that into the other equation. -7x + 8y = -5 -7(3y - 3) + 8y = -5 -21y + 21 + 8y = -5 -21y + 8y = -5 - 21 - 13y = - 26 y = -26/-13 y = 2 now sub 2 in for y in either equation to find x -2x + 6y = 6 -2x + 6(2) = 6 -2x + 12 = 6 -2x = 6 - 12 -2x = -6 x = 3 check.. -2x + 6y = 6 -2(3) + 6(2) = 6 -6 + 12 = 6 6 = 6 (correct) x = 3 and y = 2

OpenStudy (anonymous):

I don't want the answer.. I just wanna be talked through how to solve it but I'm so lost..

OpenStudy (dbzfan836):

get geogebra

OpenStudy (dbzfan836):

answer: [3,2]

OpenStudy (dbzfan836):

geogebra is a free graphin software

OpenStudy (dbzfan836):

It makes life sooo much easier